1, The following data represent the muzzle velocity​ (in feet per​ second) of rounds fired from a​ 155-mm gun. For each​ round, two measurements of the velocity were recorded using two different measuring​ devices, resulting in the following data. Complete parts​ (a) through​ (d) below.

Observation A B A-B
1 791.4 795.9 -4.5
2 791.8 791.7 0.1
3 794.1 803.2 -9.1
4 791.9 787.6 4.3
5 791.8 792.6 -0.8
6 794.8 791.6 3.2

​(c) Construct a​ 99% confidence interval about the population mean difference. Compute the difference as device A minus device B. Interpret your results.

The lower bound is

The upper bound is

​(Round to two decimal places as​ needed.)
Answer #1

= (-4.5 + 0.1 + (-9.1) + 4.3 + (-0.8) + 3.2)/6 = -1.13

sd = sqrt(((-4.5 + 1.13)^2 + (0.1 + 1.13)^2 + (-9.1 + 1.13)^2 + (4.3 + 1.13)^2 + (-0.8 + 1.13)^2 + (3.2 + 1.13)^2)/5) = 4.995

df = 6 - 1 = 5

At 99% confidence level, the critical value is t* = 4.033

The 99% confidence interval is